a ball is thrown upward with an initial velocity of 14m/s. using the approximate value of g=10m/s^2, what are the magnitude and direction of the balls velocity.

a) 1 second after it is thrown
b) 2 second after it is thrown
1
Expert's answer
2012-10-08T10:17:46-0400
The equation for velocity is

v=v0 - gt
where v0=14m/s, g=10m/s^2.
in 1 second:
v=14-10=4m/s
it is positive so direction is upwards
in 2 seconds:
v=14-20=-6m/s
it is negative so direction is downwards